# Unified Optimization of Source Weights and Transfer Quantities in Multi-Source Transfer Learning: An Asymptotic Framework

### v1

@@ III-B Problem Formulation
III-C Asymptotic Normality of the MLE
When attempting to recover the true parameter vector $\underline{\theta}^{*}$ from i.i.d observations generated according to the distribution $P_{X;\underline{\theta}^{*}}$, we let $\mathcal{D}$ denote a collection of $n$ such i.i.d. samples. The maximum likelihood estimator (MLE) is then defined as the maximizer of the empirical log-likelihood:ă
Intuitively, the Fisher information matrix quantifies the sensitivity of the likelihood function with respect to the parameter $\underline{\theta}$. A larger Fisher information implies that small changes in $\underline{\theta}$ cause more pronounced variations in the log-likelihood, meaning the parameter is easier to estimate accurately from data. This notion is also closely related to the Cramér–Rao lower bound, which states that the covariance matrix of any unbiased estimator cannot be smaller (in the positive semi-definite sense) than the inverse of the Fisher information. Therefore, the Fisher information directly characterizes the fundamental limit of parameter estimation accuracy.
IV main result
In this section, for the optimization problem in (III-B), we will first present the theoretical result and the computation method for the optimal transfer quantity and the optimal source weight in the single-source setting in Section IV-A. In the following, we extend the analysis to the multi-source setting in Section IV-B. Moreover, We introduce practical algorithms derived from our theory for multi-source transfer learning and multi-task learning in Section IV-C.
@@ IV-A Single-Source Transfer Learning
In this subsection, we first analyze the case where the parameter is one-dimensional in Theorem 2. We then extend the method to the high-dimensional setting in Proposition 3. Throughout this paper, we denote scalar-valued parameters by $\theta$ and high-dimensional parameters by $\underline{\theta}$. In analyzing the target parameter estimation problem, directly computing the K-L measure is intractable in general. Therefore, we conduct our analysis primarily in the asymptotic regime, where the behavior of the estimator becomes more tractable. To begin with, we consider the transfer learning scenario where we have one target task $\mathcal{T}$} with $N_{0}$ training samples and one source task $\mathcal{S}_{1}$ with $N_{1}$ training samples. In this case, we aim to determine the optimal transfer quantity $n_{1}^{*}\in[1,N_{1}]$ and the optimal transfer weight $w_{1}^{*}\in[0,+\infty]$. To facilitate our mathematical derivations, we assume $N_{0}$ and $N_{1}$ are asymptotically comparable, and the distance between the parameters of the target task and source task is sufficiently small (i.e., $|\theta_{0}-\theta_{1}|=O(\frac{1}{\sqrt{N_{0}}})$). Considering the similarity of low-level features among tasks of the same type, this assumption is made without loss of generality [40, 59]. Furthermore, as demonstrated in subsequent analysis, our conclusions remain valid even in extreme cases where the distance between parameters is large.
@@
We now present some observations regarding the optimal weight in (10). We observe that as the distance between $\theta_{0}$ and $\theta_{1}$ increases, the optimal weight decreases. This shows that the weighting strategy reduces the influence of sources with substantial discrepancies from the target. Moreover, as the quantity of source samples $N_{1}$ increases, the optimal weight also decreases. This illustrates that the derived weight adaptively controls the trade-off between the target and source information, preventing the source from dominating even when it is abundant. In Fig. 1, we provide an intuitive explanation for the optimal $w_{1}^{*}$.
Figure 1: The three circles represent the errors corresponding to transfer weight $w_{1}=0$, $w_{1}^{*}$, and $w_{1}>w_{1}^{*}$. The distance from each circle’s center to $\theta_{0}$ represents the bias of estimation, and the radius represents the variance. As the transfer weight of source increases, the bias term increases, while the variance term decreases for $w_{1}\in[0,1]$ and increases for $w_{1}\in(1,+\infty)$. . The optimal $w_{1}^{*}$ achieves the best trade-off between them.
In the following, we extend the result of the Theorem 2 to the high-dimensional setting.
@@
(proved in Appendix B) In the single-source setting with $d$-dimensional models $P_{X;{\underline{\theta}}_{0}}$ and $P_{X;{\underline{\theta}}_{1}}$, we assume that ${\underline{\theta}}_{0},{\underline{\theta}}_{1}\in\mathbb{R}^{d}$ and ${\underline{\theta}}_{0}-{\underline{\theta}}_{1}||=O(\frac{1}{\sqrt{N_{0}}})$. Then, the K-L measure takes the following form:
@@
Here, $t$ is a scalar ; $J({\underline{\theta}}_{0})\in\mathbb{R}^{d\times d}$ denotes the Fisher information matrix evaluated at ${\underline{\theta}}_{0}$, and $({\underline{\theta}}_{1}-{\underline{\theta}}_{0})\in\mathbb{R}^{d}$ is the coordinate-wise difference between the two parameter vectors. Finally, the optimal transfer quantity is the same form as (9) and the optimal transfer weight $w_{1}^{*}$ is the same form as (10).
Compared with Theorem 2, Proposition 3 retains an analogous structural form in its mathematical expression. This structural resemblance enables us to follow a conceptually similar analytical route to determine the corresponding optimal quantity of knowledge transfer. It is worth emphasizing that the appearance of the Fisher information matrix $J$ in the $t$-term is a direct consequence of the theoretical derivation, rather than a modeling choice or an explicitly imposed mechanism. The resulting expression admits a clear interpretation: the Fisher information matrix implicitly induces a dimension-weighted aggregation over parameter directions, where the contribution of each direction is determined by the sensitivity of the likelihood. Parameter directions along which the likelihood is more sensitive to perturbations contribute more prominently to the derived measure, while directions with lower sensitivity have a reduced influence. This interpretation clarifies that the role of $J$ is inherent to the statistical structure of the estimation problem, rather than being introduced by design.
@@
The expression in Theorem 2 and Proposition 3 coincides with the results obtained by [47]. The (23) can be transformed to
@@
This formulation exhibits a notable similarity to the conclusion presented in [47].
@@ IV-B Multi-Source Transfer Learning
In this subsection, we first present the expression of the K-L measure in the multi-source setting, as stated in Theorem 5. By minimizing this expression, we derive a method to compute the optimal source weights, and present practical algorithms for real-world multi-source transfer learning and multi-task learning scenarios in Algorithm 1 and Algorithm 2.
@@
(proved in Appendix C) In the multi-source setting with $d$-dimensional models $P_{X;\underline{\theta}_{0}}$ for the target task and $P_{X;\theta_{i}}$, $i\in[1,K]$ for the source tasks, we assume that $\underline{\theta}_{0},\underline{\theta}_{1}\dots\underline{\theta}_{K}\in\mathbb{R}^{d}$ and $\underline{\theta}_{0}-\underline{\theta}_{i}||=O(\frac{1}{\sqrt{N_{0}}})$. Then, the K-L measure is given by
@@
For optimal transfer quantities, by minimizing the (16), we obtain that maximizing the transfer quantities is optimal, i.e., $n_{i}^{*}=N_{i}$. Moreover, we propose a numerical procedure to compute the optimal source weights, which will be elaborated in the paragraph following the proof.
@@
This problem requires optimizing the objective function over two variables: a scalar variable $s$ representing the total weighted transfer quantity, and a vector variable $\underline{\alpha}$ representing the proportion of each source domain in $s$. First, we compute the optimal $\underline{\alpha}^{\prime}$ under the constraint $\mathcal{A}$ to minimize $t$, and get the optimal ${{\alpha}^{*}_{i}},i\in[1,K]$, which is a $K\times K$ non-negative quadratic programming problem with respect to $\underline{\alpha}$
@@
$\underline{\alpha}*=\arg\min\limits_{\underline{\alpha}\in\mathcal{A}}\underline{\alpha}^{T}\frac{\left(\operatorname{diag}\left(\frac{d}{N_{1}},\dots,\frac{d}{N_{K}}\right)+\Theta^{T}{J}(\underline{\theta}_{0})\Theta\right)}{d}\underline{\alpha}.$
@@
In the numerator of the quadratic coefficient matrix for this optimization problem, the first component is a diagonal matrix and is therefore strictly positive definite. The second component, $\Theta^{\top}J(\underline{\theta}_{0})\Theta$, is positive semi-definite because the Fisher information matrix $J(\underline{\theta}_{0})$ is itself positive semi-definite. By the closure property of positive semi-definite matrices under addition, the resulting quadratic coefficient matrix is also positive semi-definite. This property guarantees that the optimization problem admits at least one global optimal solution. By this procedure, we obtain $t^{*}$ and $\alpha^{*}$ that minimize $t$. Then we get the optimal value of s{s}, which is always ${s^{*}}=\frac{1}{t^{*}}$. Finally, we use optimal $s^{*}$} and ${\alpha}^{*}_{i}$ to get the optimal weights by $w^{*}_{i}=\frac{s^{*}{\alpha}^{*}_{i}}{N_{i}}$.
@@
The time complexity of this optimization method depends solely on the optimization of $\alpha$, which corresponds to a $K\times K$ non-negative quadratic programming problem, whose complexity is $\mathcal{O}(K^{3})$. In most standard transfer learning benchmarks, the number of available source domains is modest (typically $K\leq 10$), rendering the $\mathcal{O}(K^{3})$ computational complexity manageable for real-world applications. For cases where the source domain set is substantially larger, one feasible strategy to preserve scalability is to perform a preliminary clustering or aggregation of domains with high similarity before executing our algorithm on the resulting reduced set.
Remark 6 (Explanation of Using All Source Samples).
@@
The finding that maximizing the transfer quantity is optimal is non-trivial and theoretically insightful. Conventional wisdom in transfer learning holds that excessive use of source data without controlling domain discrepancy can amplify bias and cause negative transfer [60]. This perspective has motivated many approaches that either down-weight or selectively discard source samples to mitigate distribution mismatch [17, 59]. On the other hand, some methods by default exploit all source data for joint training, often relying on heuristic weighting. Our framework provides a new theoretical perspective by jointly optimizing transfer quantity and source weights. Through asymptotic analysis, we prove that when weights are allowed to adjust, the optimal solution always employs all available source samples. The key insight is that enlarging the effective sample size strictly reduces variance, while the weighting mechanism simultaneously suppresses the bias introduced by heterogeneous domains. This dual effect reconciles the seemingly conflicting intuitions in the literature: discarding data may reduce bias but sacrifices variance reduction, whereas full data utilization achieves the best trade-off once proper weighting is in place. Consequently, our analysis offers a principled justification for exploiting all source data in multi-source transfer learning.
@@
IV-C Practical Algorithms
In addition, we propose practical algorithms of UOWQ for multi-source transfer learning ă in Algorithm 1 and for multi-task learning in Algorithm 2. In the following, we focus on Algorithm 1 to illustrate the proposed approach, whose overall workflow is shown in Fig. 2. In UOWQ, when computing the optimal source weights based on Theorem 5, we use the parameters of pretrained source models to replace $\underline{\theta}_{1},\cdots,\underline{\theta}_{K}$. Considering that the source model can be trained using sufficient labeled data, it is reasonable to use the learned parameters as a good approximation of the true underlying parameters.
In contrast, the number of target data in transfer learning is often insufficient, so it is difficult to accurately estimate the true parameter $\underline{\theta}_{0}$ - the parameter of the target task model - using only the target data. Therefore, as shown in lines 5-12 of Algorithm 1, we adopt a dynamic weight update strategy. Specifically, in the first epoch, we train a $\underline{\theta}_{0}$ using only the target data. This $\underline{\theta}_{0}$ is then used, along with Theorem 5, to determine the optimal source weight of each source task. Finally, we continue training $\underline{\theta}_{0}$ with new weights, and this procedure is repeated in each subsequent epoch to iteratively update the weights. In summary, during the first epoch of the algorithm, a small set of target samples is utilized to initialize both the model parameters and the Fisher information matrix. In subsequent epochs, additional source samples are incorporated to refine this estimation, with the enlarged sample set enhancing the stability. In particular, the loss function $\ell$ is the negative log-likelihood. Moreover, in line 9, we compute the matrix $J$ using the widely adopted empirical Fisher approach in deep learning [35, 37].
V Experiments
In this section, we evaluated the practical algorithms of UOWQ in real-world datasets. In particular, we evaluated Algorithm 1 in multi-source transfer learning setting and Algorithm 2 in multi-task learning setting.
V-A Experimental Settings

### v2

@@ III-C Problem Formulation
For theoretical analysis, we assume a compact parameter space $\Theta$ and standard regularity conditions on $P_{X;\underline{\theta}}$, including differentiability, parameter-independent support, and finite, positive Fisher information.
IV Theoretical Characterization of Optimal Source Weights and Transfer Quantities
In this section, we present the main theoretical results of UOWQ. We first characterize the optimal transfer quantity and the optimal source weight in the single-source setting, deriving closed-form expressions in Section IV-A. We then extend the analysis to the multi-source setting, where the joint optimization problem admits a convex formulation and a tractable numerical solution in Section IV-B.
@@ IV-A Single-Source Transfer Learning
In this subsection, we first analyze the case where the parameter is one-dimensional in Theorem 2. We then extend the method to the high-dimensional setting in Proposition 3. Throughout this paper, we denote scalar-valued parameters by $\theta$ and high-dimensional parameters by $\underline{\theta}$. In analyzing the target parameter estimation problem, directly computing the K-L measure is intractable in general. Therefore, we conduct our analysis primarily in the asymptotic regime, where the behavior of the estimator becomes more tractable. To begin with, we consider the transfer learning scenario where we have one target task $\mathcal{T}$} with $N_{0}$ training samples and one source task $\mathcal{S}_{1}$ with $N_{1}$ training samples. In this case, we aim to determine the optimal transfer quantity $n_{1}^{*}\in[1,N_{1}]$ and the optimal transfer weight $w_{1}^{*}\in[0,+\infty]$. To facilitate our mathematical derivations, we assume $N_{0}$ and $N_{1}$ are asymptotically comparable, and the distance between the parameters of the target task and source task is sufficiently small (i.e., $|\theta_{0}-\theta_{1}|=O(\frac{1}{\sqrt{N_{0}}})$). Considering the similarity of low-level features among tasks of the same type [40], this assumption is made without loss of generality, and similar assumptions have also been adopted in the prior literature [60].
@@
In Fig. 1, we provide an intuitive explanation for the existence of the optimal weight $w_{1}^{*}$ defined in (10). We now discuss several properties of the optimal weight $w_{1}^{*}$. We observe that as the distance between $\theta_{0}$ and $\theta_{1}$ increases, $t$ increases and $w_{1}^{*}$ decreases accordingly. In particular, as $|\theta_{0}-\theta_{1}|\to 0$, $w_{1}^{*}$ approaches one, whereas as the discrepancy becomes sufficiently large, $w_{1}^{*}$ approaches zero. This shows that the weighting strategy reduces the influence of the source domain when it exhibits substantial discrepancies from the target. Moreover, as the quantity of source samples $N_{1}$ increases, $w_{1}^{*}$ also decreases. This illustrates that the derived weight adaptively controls the trade-off between the target and source information, preventing the source from dominating even when it is abundant.
Figure 1: The three circles represent the errors corresponding to transfer weight $w_{1}=0$, $w_{1}=w_{1}^{*}$, and $w_{1}>w_{1}^{*}$. The distance from each circle’s center to $\theta_{0}$ represents the bias of estimation, and the radius represents the variance. As the transfer weight of source increases, the bias term increases, while the variance term decreases for $w_{1}\in[0,1]$ and increases for $w_{1}\in(1,+\infty)$. The optimal $w_{1}^{*}$ achieves the best trade-off between them.
In the following, we extend the result of Theorem 2 to the high-dimensional setting.
@@
(proved in the Appendix) In the single-source setting with $d$-dimensional models $P_{X;{\underline{\theta}}_{0}}$ and $P_{X;{\underline{\theta}}_{1}}$, we assume that ${\underline{\theta}}_{0},{\underline{\theta}}_{1}\in\mathbb{R}^{d}$ and ${\underline{\theta}}_{0}-{\underline{\theta}}_{1}||=O(\frac{1}{\sqrt{N_{0}}})$. Then, the K-L measure takes the following form:
@@
Here, $t$ is a scalar, $J({\underline{\theta}}_{0})\in\mathbb{R}^{d\times d}$ denotes the Fisher information matrix evaluated at ${\underline{\theta}}_{0}$, and $({\underline{\theta}}_{1}-{\underline{\theta}}_{0})\in\mathbb{R}^{d}$ is the coordinate-wise difference between the two parameter vectors. Finally, the optimal transfer quantity is the same form as (9) and the optimal transfer weight $w_{1}^{*}$ is the same form as (10).
Compared with Theorem 2, Proposition 3 retains an analogous structural form in its mathematical expression. This structural resemblance enables us to follow a conceptually similar analytical route to determine the corresponding optimal quantity of knowledge transfer. It is worth emphasizing that the appearance of the Fisher information matrix $J$ in the $t$-term is a direct consequence of the theoretical derivation, rather than a modeling choice or an explicitly imposed mechanism. The resulting expression admits a clear interpretation: the Fisher information matrix implicitly induces a dimension-weighted aggregation over parameter directions, where the contribution of each direction is determined by the sensitivity of the likelihood. Parameter directions along which the likelihood is more sensitive to perturbations contribute more prominently to the derived measure, while directions with lower sensitivity have a reduced influence.
@@
The expression in Theorem 2 and Proposition 3 coincides with the results obtained by [47]. Equation (5) can be transformed to
@@
This formulation exhibits a notable similarity to the conclusions in [47]. In particular, both frameworks contain terms that scale with the inverse of the transfer quantities, highlighting the importance of accounting for sample size when determining the transfer weights. However, our framework differs from their work not only in the weighting mechanism but also in the treatment of model discrepancy and dimensionality. Specifically, Tong et al. [47] adopt a model-based weighting strategy, whereas our approach employs a sample-based weighting scheme. Their analysis quantifies task similarity via a $\chi^{2}$-distance, while the discrepancy term $t$ in our formulation measures model distance directly in the parameter space and incorporates model complexity through Fisher information. Moreover, parameter dimensionality enters the weighting mechanism differently in the two approaches.
@@ IV-B Multi-Source Transfer Learning
In this subsection, we first present the expression of the K-L measure in the multi-source setting, as stated in Theorem 5. By minimizing this expression, we derive a method to compute the optimal source weights and analyze its computational complexity. Moreover, we provide an interpretation of the theoretical result that using all available samples is optimal when the weights are adjustable.
@@
(proved in the Appendix) In the multi-source setting with $d$-dimensional models $P_{X;\underline{\theta}_{0}}$ for the target task and $P_{X;\theta_{i}}$, $i\in[1,K]$ for the source tasks, we assume that $\underline{\theta}_{0},\underline{\theta}_{1}\dots\underline{\theta}_{K}\in\mathbb{R}^{d}$ and $\underline{\theta}_{0}-\underline{\theta}_{i}||=O(\frac{1}{\sqrt{N_{0}}})$. Then, the K-L measure is given by
@@
For optimal transfer quantities, by minimizing (16), we obtain that maximizing the transfer quantities is optimal, i.e., $n_{i}^{*}=N_{i}$. Moreover, we propose a method to compute the optimal source weights, which will be elaborated in the following paragraph.
@@
This problem requires optimizing the objective function over two variables: a scalar variable $s$ representing the total weighted transfer quantity, and a vector variable $\underline{\alpha}$ representing the proportion of each source domain in $s$. First, we compute the optimal $\underline{\alpha}^{*}=\left[\alpha_{1}^{*},\dots,\alpha_{K}^{*}\right]^{T}$ under the constraint $\mathcal{A}$ to minimize $t$, which is a $K\times K$ non-negative quadratic programming problem with respect to $\underline{\alpha}$, i.e.,
@@
$\underline{\alpha}^{*}=\arg\min\limits_{\underline{\alpha}\in\mathcal{A}}\underline{\alpha}^{T}\frac{\left(\operatorname{diag}\left(\frac{d}{N_{1}},\dots,\frac{d}{N_{K}}\right)+\Theta^{T}{J}(\underline{\theta}_{0})\Theta\right)}{d}\underline{\alpha}.$ (23)
@@
In the quadratic coefficient matrix of this optimization problem, the first component, $\operatorname{diag}\left(\frac{d}{N_{1}},\dots,\frac{d}{N_{K}}\right)$, is diagonal and therefore strictly positive definite. The second component, $\Theta^{\top}J(\underline{\theta}_{0})\Theta$, is positive semi-definite since the Fisher information matrix $J(\underline{\theta}_{0})$ is positive definite by the assumptions in Section III. Consequently, the sum of these two components is positive definite, which guarantees that the optimization problem is convex and admits a unique global optimum. By this procedure, we obtain $t^{*}$ and $\alpha^{*}$ that minimize $t$. Then we get the optimal value of s{s}, which is always
${s^{*}}=\frac{1}{t^{*}}.$ (24)
Finally, we use optimal $s^{*}$} and ${\alpha}^{*}_{i}$ to get the optimal weights by ă
$w^{*}_{i}=\frac{s^{*}{\alpha}^{*}_{i}}{N_{i}}.$ (25)
@@
The time complexity of this optimization is dominated by solving a $K\times K$ quadratic programming problem over $\alpha$, which can be solved in $\mathcal{O}(K^{3})$ time using a primal-dual interior-point solver. In most standard transfer learning benchmarks, the number of available source domains is modest (typically $K\leq 10$), rendering the $\mathcal{O}(K^{3})$ computational complexity manageable for real-world applications. For cases where the number of available source domains is substantially larger, one feasible strategy to preserve scalability is to perform a preliminary clustering or aggregation of domains with high similarity before executing our algorithm.
@@
Moreover, the finding that maximizing the transfer quantity is optimal is non-trivial and theoretically insightful. Conventional wisdom in transfer learning holds that excessive use of source data without controlling domain discrepancy can amplify bias and cause negative transfer [61]. This perspective has motivated many approaches that either down-weight or selectively discard source samples to mitigate distribution mismatch [18, 23, 60]. On the other hand, some methods by default exploit all source data for joint training, often relying on heuristic weighting. Our framework provides a new theoretical perspective by jointly optimizing transfer quantity and source weights. Through asymptotic analysis, we prove that when weights are allowed to adjust, the optimal solution always employs all available source samples. The key insight is that enlarging the effective sample size strictly reduces variance, while the weighting mechanism simultaneously suppresses the bias introduced by heterogeneous domains. This dual effect reconciles prior intuitions: discarding data may reduce bias but weakens variance reduction, while our analysis shows that full data utilization with proper weighting achieves a better bias–variance trade-off. Consequently, our analysis offers a principled justification for exploiting all source data in multi-source transfer learning.
@@
V Practical Algorithms
In this section, we introduce practical algorithms derived from our theory for multi-source transfer learning and multi-task learning. Specifically, we propose practical algorithms of UOWQ for multi-source transfer learning in Algorithm 1 and for multi-task learning in Algorithm 2.
Firstly, we focus on Algorithm 1 to illustrate the proposed approach, whose overall workflow is shown in Fig. 2. In UOWQ, when computing the optimal source weights based on Theorem 5, we use the parameters of pretrained source models to replace $\underline{\theta}_{1},\cdots,\underline{\theta}_{K}$. Considering that the source model can be trained using sufficient labeled data, it is reasonable to use the learned parameters as a good approximation of the true underlying parameters. In contrast, the amount of target data in transfer learning is often insufficient, so it is difficult to accurately estimate the true parameter $\underline{\theta}_{0}$ - the parameter of the target task model - using only the target data. Therefore, as shown in lines 5-12 of Algorithm 1, we adopt a dynamic weight update strategy. Specifically, in the first epoch, we train $\underline{\theta}_{0}$ with the source weights initialized to zero, which is equivalent to using only the target data. This $\underline{\theta}_{0}$ is then used, along with Theorem 5, to determine the optimal source weight for each source task. Thereafter, training proceeds in an alternating manner: given the current weights, we update $\underline{\theta}_{0}$, and given the updated model, we recompute the weights. This process is repeated at each epoch, enabling iterative refinement of both the model and the weights. Such an iterative scheme allows adaptive weight adjustment, resulting in more stable optimization and improved transfer performance. In particular, the loss function $\ell$ is the negative log-likelihood. Moreover, in line 9, we compute the matrix $J$ using the widely adopted empirical Fisher approach in deep learning [36, 38].
We further extend the proposed framework to the multi-task learning setting, where each task simultaneously serves as both a target and a source for the others, and the objective is to learn task-specific models for all tasks. Suppose we are given $K$ tasks $\{S_{k}\}_{k=1}^{K}$, where each task $S_{k}$ is associated with dataset $D_{S_{k}}$ and parameter $\theta_{k}$. When optimizing task $k$, we treat $S_{k}$ as the target task and the remaining tasks $\{S_{k^{\prime}}\}_{k^{\prime}\neq k}$ as auxiliary sources, and we construct its training set by combining all samples from the target task with the weighted samples from the remaining tasks, where the weights $w_{k^{\prime}}^{(k)}$ are computed by UOWQ. The resulting problem becomes a joint optimization over all task parameters $\{\theta_{k}\}_{k=1}^{K}$, leading to an alternating optimization procedure between tasks, summarized in Algorithm 2.
VI Experiments
VI-A Experimental Settings
In the experiments, we evaluated the practical algorithms of UOWQ in real-world datasets. In particular, we evaluated Algorithm 1 in the multi-source transfer learning setting and Algorithm 2 in the multi-task learning setting.